Hi.
I'm checking to see if I got this problem correct.
Solve the equation.
3y/4-13=-y/3
thanks
H.C.

Answer by edjones(8007)

3y/4 - 13=-y/3
9y-156=-4y
13y=156
y=12
.
Ed